Two wires separated by 10 cm carry a current in the plane of the page as shown below. Determine the magnitude and direction (draw or describe it) of the net magnetic field, due to these two currents, at point ? located 5 cm to the right of the right-most wire.

10 cm 12 = 20 A 5 cm 11 = 30 A P
Answer #1

From right hand palm rule the magnetic field due to I1 is into the page and due to I2 is out of the page

Now total magnetic field will be

B1-B2

B1= uI1/2πR1 = 2*10^-7*30/0.15

B2 = 2*10^-7*20/0.05

B1-B2 =-4*10^-5T into the page

Or B= 4*10^-5 T out of page